# Generating Realistic Arm Movements in Reinforcement Learning: A Quantitative Comparison of Reward Terms and Task Requirements

### v1

@@ Abstract
The mimicking of human-like arm movement characteristics involves the consideration of three factors during control policy synthesis: (a) chosen task requirements, (b) inclusion of noise during movement execution and (c) chosen optimality principles. Previous studies showed that when considering these factors (a-c) individually, it is possible to synthesize arm movements that either kinematically match the experimental data or reproduce the stereotypical triphasic muscle activation pattern. However, to date no quantitative comparison has been made on how realistic the arm movement generated by each factor is; as well as whether a partial or total combination of all factors results in arm movements with human-like kinematic characteristics and a triphasic muscle pattern. To investigate this, we used reinforcement learning to learn a control policy for a musculoskeletal arm model, aiming to discern which combination of factors (a-c) results in realistic arm movements according to four frequently reported stereotypical characteristics. Our findings indicate that incorporating velocity and acceleration requirements into the reaching task, employing reward terms that encourage minimization of mechanical work, hand jerk, and control effort, along with the inclusion of noise during movement, leads to the emergence of realistic human arm movements in reinforcement learning. We expect that the gained insights will help in the future to better predict desired arm movements and corrective forces in wearable assistive devices.
@@ I Introduction
Arm-reaching movements exhibit highly stereotypical kinematics and temporal characteristics. Important characteristics that have been documented in literature are: (i) roughly straight hand trajectories, (ii) bell-shaped tangential velocity profiles [5, 6], (iii) triphasic muscle activation pattern [7, 8], i.e., the alternating activation of agonist and antagonist muscles, and (iv) linear relationship between movement time (MT) and index of difficulty (ID) (a.k.a Fitts’s law) [9]. Several optimality principles have been proposed for deterministic prediction of arm-reaching movements, such as minimal work, jerk or muscular effort [10, 11]. Flash et al. [11] found that minimization of hand jerk predicts characteristics (i&ii) on point-to-point movements. Wochner et al. [12] indicated that minimization of mechanical work, jerk, and muscle stimulation command (effort) predicts characteristics (i&ii) on point-to-manifold movements. Finally, Ueyama et al. [13] demonstrated that minimization of control effort and consideration of position, velocity, and force requirements in the reaching task predict characteristics (i-iii). As a stochastic approach, Fischer et al. [4] applied constant and signal-dependent noise of muscle stimulation amplitude. Combined with minimization of movement time they were able to reproduce characteristics (i&ii&iv) on point-to-point movements. To our knowledge, there is no simulation approach that investigated all four characteristics.
More precisely, three factors influence the resulting behaviour of the control policy to generate human characteristics of arm movement: (a) the chosen task requirements, (b) inclusion of noise during movement execution and (c) the chosen optimality principles. Some of these factors have been evaluated based on their ability to generate kinematic characteristics that match experimental data, while others evaluated the emergence of the triphasic muscle activation pattern. However, to date no quantitative comparison has been conducted on how realistic the arm movement generated by each factor is; as well as whether a partial or total combination of all factors result in arm movements with human-like kinematic features and muscle activation pattern.
@@ II-A Muscle stimulation commands
The RL agent utilizes Maximum a Posteriori Optimization (MPO) [14] combined with DEP-RL [15] for exploration; a novel approach that demonstrated robust performance controlling musculoskeletal systems. The MPO implementation follows the default settings provided by the TonicRL library [16]111except for the following parameters: batch size with $256$, batch iteration with $30$, steps before batches with $51e$ 5 and steps between batches with $50$., and the DEP-RL configuration mirrors the hyperparameters outlined for the same arm model in [15].
@@ II-B Simulation of human arm dynamics
The physics engine MuJoCo [18] simulates the muscle activation dynamics resulting in the generation of muscle forces that drive the arm movements. In MuJoCo, a musculoskeletal model of the human arm with two degrees of freedom and six actuated muscles is available [18]. The original model was modified to generate arm movements in the sagittal plane (considering gravity). The position error is calculated between the tip of the forearm and the desired position (reaching goal). The RL environment considers the same initial conditions of the arm model for all episodes: $0^{\circ}$ for the shoulder angle, $90^{\circ}$ for the elbow angle, zero joint velocities and zero muscle activation level. The environment observation comprises Cartesian states222position, velocity and acceleration of the hand, i.e., tip of the forearm. and joint states333position, velocity, acceleration and jerk of each arm joint, muscle states444muscle activity, muscle forces, muscle lengths and muscle velocities, mechanical work, hand jerk and the goal position. The agent’s policy network generates control commands every $10$ ms while the MuJoCo physics engine updates the arm model every $2$ ms and uses the same control command for five consecutive time steps.
@@ II-C Reward formulation
Generally, arm pointing movements are executed quickly. Thus, previous research used a constant negative reward for every step transition until the force requirement is met [4]. In addition, we incorporate velocity and acceleration requirements into the reaching task. Therefore, $r_{\mathrm{sparse}}$ depends on fulfillment of goal tolerance in position and the additional kinematics requirements
$r_{\mathrm{sparse}}=\left\{\begin{array}[]{@{}ll@{}}0,&\text{if}\ ||p_{\mathrm{hand}}-p_{\mathrm{goal}}||\leq p_{\mathrm{tol}}\And\mathrm{t}_{\mathrm{req}}=\text{true}\\ -1,&\text{otherwise},\end{array}\right.$
@@
$\mathrm{t}_{\mathrm{req}}=\left\{\begin{array}[]{@{}ll@{}}\text{true},&\text{pos}\\ ||v||\leq v_{\mathrm{tol}},&\text{pos-vel}\\ ||v||\leq v_{\mathrm{tol}}\And||a||\leq a_{\mathrm{tol}},&\text{pos-vel-acc}\\ \end{array}\right.$
where $v,a$ are hand velocity and acceleration, $v_{\mathrm{tol}},a_{\mathrm{tol}}$ are tolerance for velocity and acceleration. We choose $v_{\mathrm{tol}}$, $a_{\mathrm{tol}}$ to be $10\%$ of the maximum values observed solely under position task requirement: $v_{\mathrm{tol}}=20$ $\frac{\mathrm{cm}}{\mathrm{s}}$ and $a_{\mathrm{tol}}=10$ $\frac{\mathrm{cm}}{\mathrm{s}^{2}}$.
@@
Only considering movement time minimization will generate bang-bang control solutions with asymmetric velocity profiles [19]. For this reason, previous work considered optimal behavior based on minimization of mechanical work, hand jerk and muscle stimulation command [12]. In this way, $r_{\mathrm{optimal}}$ considers three optimality principles as
@@
where $c_{3,4,5}$ set priority between optimality principles, $r_{\mathrm{effort}}$ is computed as mean value of muscle stimulation commands, $r_{\mathrm{jerk}}$ is estimated by finite difference computation between the current and one previous acceleration values, instantaneous work (power) $r_{\mathrm{work}}$ is computed as $|\dot{\phi}_{1}\tau_{1}|+|\dot{\phi}_{2}\tau_{2}|$, where $\dot{\phi}_{1,2}$ represents the angular velocity of shoulder and elbow, and $\tau_{1,2}$ indicates the torque of shoulder and elbow. We normalize each optimality principle by its observed maximum value: $r_{\mathrm{jerk}}=1000$ $\frac{\text{m}}{\text{s}^{3}}$ and $r_{\mathrm{work}}=100$ J. In pre-tests, smooth muscle profiles were achieved with the following coefficients: $c_{3}=1$, $c_{4}=8$, and $c_{5}=1$.
@@ II-D Metrics for goal-oriented movements
    Triphasic muscle pattern ($u_{\text{triphasic}}$): This metric analyzes the muscle activation pattern of each agonist-antagonist muscle pair in the arm. The aim is to capture if an antagonistic pair changes operation mode, e.g., if in the beginning elbow flexor is actively flexing the elbow and then, elbow extensor activity rises and elbow flexor activity falls to decelerate the movement, this is considered a second phase. We quantify this by evaluating if muscle activation slopes exchange directions and by the threshold $\Delta>0.25\Delta_{\mathrm{max}}\And\Delta>1.5\mathrm{e}-3$ of difference between them. If this is the case, it is considered a new phase of muscle activation. The metric verifies if the reported triphasic pattern in the literature [13] occurs in a muscle pair, assigning a score of $1$ if true and 00 otherwise.
@@ III-B Bell-shaped velocity profile ($v_{\mathrm{bell}}$)
[…]
The best performance in terms of triphasic muscle pattern $u_{\text{triphasic}}$ across the majority of difficulty indices ($\mathrm{ID}=2$ to $5$) for all models (Baseline, Execution noise, Optimality principles and Hybrid) is linked to pos-vel and pos-vel-acc task requirements (Tab. I). The muscle patterns for $\mathrm{ID}=5$ of hybrid model for each task requirement are shown in Figure 4. The figure illustrates that both pos-vel and pos-vel-acc task requirements give rise to a triphasic muscle pattern in the elbow muscle pair, whereas only position task requirement (pos) results in a biphasic pattern in the three muscle pairs. Furthermore, the figure displays two triphasic muscle patterns for the pos-vel-acc task requirement. Similarly, Figure 5 illustrates that third muscle duration increases with larger index of difficulty (ID).
@@ III-D Fitts’s law ($R_{F}$)
The models Execution noise, Optimality principles and Hybrid obtain their highest correlation coefficient $R_{F}$ when incorporating velocity requirement into the main task (pos-vel) (Tab. I). In contrast, the Baseline model attains its highest correlation coefficient $R_{F}$ when employing only the position task requirement. The optimality principles model demonstrate slightly superior performance in $R_{F}$ compared to Baseline, Execution noise and Hybrid models. The linear relationship between Movement Time (MT) and Index of Difficulty (ID) is graphically illustrated in Figure 6 for the optimality principle model considering the three possible task requirements. Additionally, the figure illustrates the increase in movement time variance as more kinematic requirements are incorporated into the main task. It is crucial to emphasize that all combinations generate a robust linear relationship, with correlation coefficients $R_{F}>0.9$. Consequently, although certain combinations exhibit higher correlation coefficients than others, it is implausible to indicate which combination yields the most realistic arm trajectory based solely on the $R_{F}$.
@@ IV Discussion
Through the systematic combination of factors (a-c), we identify three key considerations for generating human-like characteristics in point-to-point arm movements. First, including both velocity and acceleration as task requirement ((a), pos-vel-acc), results in good or excellent values across all metrics and in the majority of difficulty indices, regardless of the model. Second, using noise b) during movement execution, in combination with reward terms c) minimizing mechanical work, hand jerk, and control effort results in the most bell-shaped velocity profile across the majority of difficulty indexes (except for the position-only task requirement). Third, a higher endpoint accuracy, i.e. a larger index of difficulty (ID), results in longer duration of the third muscle phase and velocity profiles with better bell shape. According to Wierzbicka et al. [7], the role of this third phase is to regulate the braking forces to guide the hand towards the target position. Based on that, the effect of the index of difficulty (ID) could be understood as the third phase arose to achieve the high endpoint accuracy, which in turn may have smoothed the velocity profiles on the right side, improving the bell shape.
@@
While our used control approach leads to realistic arm movements with human-like characteristics, our study presents some limitations as the used arm model includes only two degrees of freedom. Furthermore, the investigated task includes only point-to-point reaching tasks, whereas more openly defined tasks such as point-to-manifold reaching might be interesting for future research, as they offer more freedom in arm movement generation. Previous studies [10, 12] have shown significance differences in the generated arm trajectories using point-to-manifold reaching that have not been observed in point-to-point movements. Moreover, complex movements in a complex arm model may further distinguish between the different combinations such that a solution for predicting realistic human arm movements with RL could aid the development and control of assistive devices.

### v2

@@ Abstract
Mimicking of human-like arm movement characteristics involves considering three factors during control policy synthesis: (a) task requirements, (b) noise during movement execution, and (c) optimality principles. Previous studies showed that when these factors (a-c) are considered individually, it is possible to synthesize arm movements that either kinematically match experimental data or reproduce the stereotypical triphasic muscle activation pattern. However, no quantitative comparison has assessed the realism of arm movements generated by each factor, nor has it been determined whether combining these factors results in movements with human-like kinematic characteristics and the triphasic muscle pattern. To investigate this, we used reinforcement learning to learn a control policy for a musculoskeletal arm model, aiming to discern which combination of factors (a-c) results in realistic arm movements according to four frequently reported stereotypical characteristics. Our findings indicate that incorporating velocity and acceleration requirements into the reaching task, employing reward terms that minimize mechanical work, hand jerk, and control effort, along with the inclusion of noise during movement, leads to realistic human arm movements by reinforcement learning. We expect that the gained insights will help in the future to better predict desired arm movements and corrective forces in wearable assistive devices.
@@ I Introduction
Arm-reaching movements exhibit highly stereotypical kinematics and temporal characteristics. Important characteristics documented in literature are: (i) roughly straight hand trajectories, (ii) bell-shaped tangential velocity profiles [5, 6], (iii) triphasic muscle activation pattern [7, 8], i.e., the alternating activation of agonist and antagonist muscles, and (iv) linear relationship between movement time (MT) and index of difficulty (ID) (a.k.a Fitts’s law) [9]. Several optimality principles have been proposed for deterministic prediction of arm-reaching movements, such as minimal work, jerk or muscular effort [10, 11]. Flash et al. [11] found that minimization of hand jerk predicts characteristics (i&ii) in point-to-point movements. Wochner et al. [12] indicated that minimization of mechanical work, jerk, and muscle stimulation command (effort) predicts characteristics (i&ii) in point-to-manifold movements. Finally, Ueyama et al. [13] demonstrated that minimization of control effort and consideration of position, velocity, and force requirements in the reaching task predict characteristics (i-iii). As a stochastic approach, Fischer et al. [4] applied constant and signal-dependent noise of muscle stimulation amplitude. Combined with minimization of movement time they were able to reproduce characteristics (i&ii&iv) on point-to-point movements. To our knowledge, no simulation approach has investigated all four characteristics.
More precisely, three factors influence the resulting behavior of the control policy to generate human characteristics of arm movement: (a) the chosen task requirements, (b) inclusion of noise during movement execution and (c) the chosen optimality principles. Some of these factors have been evaluated based on their ability to generate kinematic characteristics that match experimental data, while others evaluated the emergence of the triphasic muscle activation pattern. However, no quantitative comparison has been conducted on the realism of the arm movement generated by each factor; as well as whether a partial or total combination of all factors results in arm movements with human-like kinematic and muscle activation pattern.
@@ II-A Muscle stimulation commands
The RL agent utilizes Maximum a Posteriori Optimization (MPO) [14] combined with DEP-RL [15] for exploration; a novel approach that demonstrated robust performance controlling musculoskeletal systems. The MPO implementation follows the default settings provided by the TonicRL library [16]111except for the following parameters: batch size with $256$, batch iteration with $30$, steps before batches with $61e$ 6 and steps between batches with $50$., and the DEP-RL configuration mirrors the hyperparameters outlined for the same arm model in [15].
@@ II-B Simulation of human arm dynamics
The physics engine MuJoCo [18] simulates the muscle activation dynamics resulting in the generation of muscle forces that drive the arm movements. In MuJoCo, a musculoskeletal model of the human arm with two degrees of freedom and six actuated muscles is available [18]. The original model was modified to generate arm movements in the sagittal plane (considering gravity). The position error is calculated between the tip of the forearm and the desired position (reaching goal). The RL environment considers the same initial conditions of the arm model for all episodes: $0^{\circ}$ for the shoulder angle, $90^{\circ}$ for the elbow angle, zero joint velocities and zero muscle activation level. The environment observation comprises Cartesian states222position, velocity and acceleration of the hand, i.e., tip of the forearm., joint states333position, velocity, acceleration and jerk of each arm joint, muscle states444muscle activity, muscle forces, muscle lengths and muscle velocities, mechanical work, hand jerk and the goal position. The agent’s policy network generates control commands every $10$ ms while the MuJoCo physics engine updates the arm model every $2$ ms and uses the same control command for five consecutive time steps.
@@ II-C Reward formulation
Previous studies have successfully generated reaching arm movements utilizing an optimal control framework [13, 12]. This methodology incorporates a terminal cost that penalizes deviations from a desired final state and an accumulated cost associated with the states and control commands during the trajectory [19]. Building upon this, the reward function consists of a sparse reward linked to the fulfillment of kinematic requirements at the end of the trajectory and a dense reward associated with executing optimal movements.
@@
Generally, arm-pointing movements are executed quickly. Consequently, previous research employed a constant negative reward for each step transition until the position requirement is met [4]. Additionally, Ueyama et al. [13] found that velocity and force requirements influence the stereotypical triphasic activation pattern. Considering these findings, we incorporate terminal velocity and acceleration (proportional to force in Cartesian space) requirements into the reaching task. Therefore, $r_{\mathrm{sparse}}$ depends on meeting the goal tolerance in position as well as the additional kinematic requirements.
$r_{\mathrm{sparse}}=\left\{\begin{array}[]{@{}ll@{}}0,&\text{if}\ \|p_{\mathrm{hand}}-p_{\mathrm{goal}}\|\leq p_{\mathrm{tol}}\And\mathrm{t}_{\mathrm{req}}=\text{true}\\ -1,&\text{otherwise},\end{array}\right.$ (2)
@@
$\mathrm{t}_{\mathrm{req}}=\left\{\begin{array}[]{@{}ll@{}}\text{true},&\text{pos}\\ \|v\|\leq v_{\mathrm{tol}},&\text{pos-vel}\\ \|v\|\leq v_{\mathrm{tol}}\And\|a\|\leq a_{\mathrm{tol}},&\text{pos-vel-acc}\\ \end{array}\right.$
where $v,a$ are hand velocity and acceleration, $v_{\mathrm{tol}},a_{\mathrm{tol}}$ are tolerance for velocity and acceleration. We choose $v_{\mathrm{tol}}$, $a_{\mathrm{tol}}$ to be $10\%$ of the maximum values observed solely under position task requirement: $v_{\mathrm{tol}}=20\,\frac{\mathrm{cm}}{\mathrm{s}}$ and $a_{\mathrm{tol}}=100\,\frac{\mathrm{cm}}{\mathrm{s}^{2}}$.
@@
Exclusively focusing on minimizing movement time will generate bang-bang control solutions with asymmetric velocity profiles [20]. Wochner et al. [12] found that bell-shaped velocity profiles emerge in point-to-manifold tasks only when optimal behavior considers the minimization of mechanical work, hand jerk, and muscle stimulation commands (related to muscular effort). As both Berret et al. [10] and Wochner et al. [12] suggested that it is crucial to consider a combination of optimality principles to tackle the redundancy problem, we therefore, consider the suggested combination $r_{\mathrm{optimal}}$ of three optimality principles as:
@@
where $c_{3,4,5}$ set priority between optimality principles, $r_{\mathrm{effort}}$ is computed as mean value of muscle stimulation commands, $r_{\mathrm{jerk}}$ is estimated by finite difference computation between the current and one previous acceleration values, instantaneous work (power) $r_{\mathrm{work}}$ is computed as $|\dot{\phi}_{1}\tau_{1}|+|\dot{\phi}_{2}\tau_{2}|$, where $\dot{\phi}_{1,2}$ represents the angular velocity of shoulder and elbow, and $\tau_{1,2}$ indicates the torque of shoulder and elbow. We normalize each optimality principle by its observed maximum value: $r_{\mathrm{jerk}}=1000\,\frac{\text{m}}{\text{s}^{3}}$ and $r_{\mathrm{work}}=100\,\text{J}$. In pre-tests, we found that smooth muscle profiles were only achieved if all the three terms are considered with the following coefficients: $c_{3}=1$, $c_{4}=8$, and $c_{5}=1$.
@@ II-D Metrics for goal-oriented movements
    Triphasic muscle pattern ($u_{\text{triphasic}}$): This metric analyzes the muscle activation pattern of each agonist-antagonist muscle pair in the arm. The aim is to capture if an antagonistic pair changes operation mode, e.g., if in the beginning elbow flexor is actively flexing the elbow and then, elbow extensor activity rises and elbow flexor activity falls to decelerate the movement, this is considered a second phase. We quantify this by evaluating if muscle activation slopes exchange directions and by the threshold $\Delta>0.25\,\Delta_{\mathrm{max}}$ and $\Delta>1.5\,\mathrm{e}-3$ of difference between them. If this is the case, it is considered a new phase of muscle activation. The metric verifies if the reported triphasic pattern in the literature [13] occurs in a muscle pair, assigning a score of $1$ if true and 00 otherwise.
@@ III-B Bell-shaped velocity profile ($v_{\mathrm{bell}}$)
III-C Triphasic muscle pattern ($u_{\text{triphasic}}$)
The best performance in terms of triphasic muscle pattern $u_{\text{triphasic}}$ across the majority of difficulty indices ($\mathrm{ID}=2$ to $5$) for all models (Baseline, Execution noise, Optimality principles and Hybrid) is linked to pos-vel and pos-vel-acc task requirements (Tab. I). The muscle patterns for $\mathrm{ID}=5$ of hybrid model for each task requirement are shown in Figure 4. The figure illustrates that both pos-vel and pos-vel-acc task requirements give rise to a triphasic muscle pattern in the elbow muscle pair, whereas only position task requirement (pos) results in a biphasic pattern in the three muscle pairs. Furthermore, the figure displays two triphasic muscle patterns for the pos-vel-acc task requirement. Similarly, Figure 5 illustrates that duration of the third muscle phase increases with larger index of difficulty (ID).
@@ III-D Fitts’s law ($R_{F}$)
The models Execution noise, Optimality principles and Hybrid obtain their highest correlation coefficient $R_{F}$ when incorporating velocity requirement into the main task (pos-vel) (Tab. I). In contrast, the Baseline model attains its highest correlation coefficient $R_{F}$ when employing only the position task requirement. The optimality principles model demonstrate slightly superior performance in $R_{F}$ compared to Baseline, Execution noise and Hybrid models. The linear relationship between Movement Time (MT) and Index of Difficulty (ID) is graphically illustrated in Figure 6 for the Optimality Principles model considering the three possible task requirements. Additionally, the figure illustrates the increase in movement time variance as more kinematic requirements are incorporated into the main task. It is crucial to emphasize that all combinations generate a robust linear relationship, with correlation coefficients $R_{F}>0.9$. Consequently, although certain combinations exhibit higher correlation coefficients than others, it is implausible to indicate which combination yields the most realistic arm trajectory based solely on the $R_{F}$.
@@ IV Discussion
Through the systematic combination of factors (a-c), we identify three key considerations for generating human-like characteristics in point-to-point arm movements. First, including both velocity and acceleration as task requirement ((a), pos-vel-acc), results in good or excellent values across all metrics and in the majority of difficulty indices, regardless of the model. Second, using noise b) during movement execution, in combination with reward terms c) minimizing mechanical work, hand jerk, and control effort results in the most bell-shaped velocity profile across the majority of difficulty indexes (except for the position-only task requirement). Third, a higher endpoint accuracy, i.e., a larger index of difficulty (ID), leads to a longer duration of the third muscle phase and velocity profiles with a better-defined bell shape. According to Wierzbicka et al. [7], the role of this third phase is to regulate the braking forces to guide the hand towards the target position. Therefore, the effect of the index of difficulty (ID) can be understood as the prolongation of the deceleration phase to achieve high endpoint accuracy, which in turn smooths the velocity profiles on the right side, enhancing the bell shape. It is noteworthy that although increasing the index of difficulty (ID) has improved the bell shape, larger values will cause the velocity profile to become more positively asymmetric [21].
@@
Although our control approach generates realistic arm movements with human-like characteristics, our study has some limitations. The arm model used incorporates only two degrees of freedom and six muscles. Consequently, our model does not fully account for the entire joint and muscle redundancy found in a real human arm. Furthermore, the investigated task includes only point-to-point reaching tasks, whereas more openly defined tasks such as point-to-manifold reaching might be interesting for future research, as they offer more freedom in arm movement generation. Previous studies [10, 12] have shown significance differences in the generated arm trajectories using point-to-manifold reaching that have not been observed in point-to-point movements. Moreover, complex movements in a complex arm model may further distinguish between the different combinations such that a solution for predicting realistic human arm movements with RL could aid the development and control of assistive devices.